# A Universal Analytic Model for Gravitational Lensing by Self-Interacting Dark Matter Halos

### v1

@@ Abstract
We present a model for analytically calculating gravitational lensing by self-interacting dark matter (SIDM) halos. Leveraging the universal behavior of SIDM halos during gravothermal evolution, we calibrate the lensing potential using a fluid simulation, normalizing the evolution time to align with established scenarios. From this potential, we derive explicit equations for the deflection angle and surface density profile, quantifying their deviations from numerical results. Our model builds on the parametric approach of ref. [1], providing refinements in the deep core-collapse regime and enabling more comprehensive lensing studies. We explore characteristic lensing features, including critical curves and caustics, for SIDM halos in isolation and within a main halo, tracking their evolution through the gravothermal phase. We also examine signatures in the self-similar regime of core collapsed halos and highlight the role of baryonic effects in realistic halos. The efficiency of our model enables large-scale lensing studies, and we make our implementation publicly available on GitHub to support further research.
@@ 1 Introduction
The paper is organized as follows. In section 2, we provide the theoretical background and describe the simulation data used for calibration. Section 3 develops an analytic model for SIDM lensing studies and evaluates its performance against numerical results. Section 4 explores the lensing signatures derived from our model, focusing on isolated cluster halos and subhalos embedded within central cluster halos. Section 5 presents a detailed analysis of lensing signatures of core collapsed halos in the self-similar regime. In section 6, we explore the impact of baryons, using a contracted density profile as an estimate. Section 7 summarizes our results, discussing potential applications of our model and directions for future development. For numerical illustrations, we assume a flat $\Lambda$CDM cosmology with $h=0.7$, $\Omega_{m}=0.3$, $\Omega_{\Lambda}=0.7$, and $w_{0}=-1$. Our calibrated models are independent of these cosmological parameters.
@@ 2 Universality and simulation data for model calibration
We present a visual representation of the universal gravothermal evolution in figure 1, which can be interpreted as a general solution for a gravothermal system: Given a halo and an SIDM model, the evolution can be derived by specifying the corresponding $\rho_{s,0}$, $r_{s,0}$, and $t_{c}$. The density profile is normalized by the NFW scale parameters at $\tau=0$, with its evolution scaled by $t_{c}$. The corresponding logarithmic density slope, $d\ln(\rho/\rho_{s,0})/d\ln(r/r_{s,0})$, is shown in the right panel. At $\tau=0$, the density profile follows the NFW form, characterized by an inner logarithmic slope of $-1$. By $\tau\approx 1.1$, the density transitions to a power-law profile, $r^{-2.2}$, featuring a core-collapsed halo in the LMFP regime [89, 40]. Dotted and dashed lines are used to track the evolution of the inner halo density, $\rho_{c}$, at $r/r_{s,0}=0.01$, and the core size, $r_{c}$, within the parametric model. Following a sharp decline at $\tau\gtrsim 0$, $\rho_{c}$ reaches its minimum near $\tau\lesssim 0.2$, after which it stabilizes into a quasi-stable core that contracts slowly. As $\tau$ approaches unity, core contraction accelerates significantly, indicating the onset of runaway collapse.
@@
Figure 2: Comparison of SIDM halo density profiles from simulations and model predictions at representative $\tau$ values. The N-body simulation results from ref. [1] were used to calibrate the $\beta 4$ model (orange), while the fluid simulation based on the program of refs. [69, 85] was used to calibrate the PSIDM model (blue) in this work. As discussed in the main text, differences between the previous and new models are minimized for $\tau<0.7$, beyond which the fluid simulation results are expected to provide greater accuracy, particularly in the inner regions.
@@ 3.1 The lensing model
[…]
@@
$a(\tau)$ $=1.113-1.262\cdot\tau^{0.1}+2.876\cdot\tau^{0.2}-3.616\cdot\tau^{0.5}+2.778\cdot\tau^{0.9}-0.4159\cdot\tau^{2}$ (3.5)
$\quad+0.01357\cdot\tau^{12}-0.03862\cdot\tau^{20}-0.0002633\cdot\tau^{87},$
$b(\tau)$ $=6.597+2.213\cdot\tau^{0.01}-9.563\cdot\tau^{0.2}+9.104\cdot\tau^{0.6}-4.013\cdot\tau^{0.9}+3.268\cdot\tau^{2}+2.076\cdot\tau^{12}$
$\quad+0.1184\cdot\tau^{51}-0.0002105\cdot\tau^{109}+4.214\cdot 10^{-7}\cdot\tau^{203},$
$c(\tau)$ $=1.793+0.5108\cdot\tau^{0.1}-1.430\cdot\tau^{0.2}+1.392\cdot\tau^{0.4}-0.7858\cdot\tau^{0.9}+0.1604\cdot\tau^{2}$
$\quad+0.03146\cdot\tau^{12}+0.0007148\cdot\tau^{67},$
$p(\tau)$ $=6.943+4.967\cdot\tau^{0.2}-12.53\cdot\tau^{0.3}+7.546\cdot\tau^{0.8}+3.212\cdot\tau^{2}+2.863\cdot\tau^{11}-0.3265\cdot\tau^{22}$
$\quad+0.3093\cdot\tau^{37},$
$s(\tau)$ $=1.825-0.5991\cdot\tau^{0.2}+0.8449\cdot\tau^{0.3}-0.4346\cdot\tau^{0.8}+0.08176\cdot\tau^{2}+0.01675\cdot\tau^{11}$
$\quad-0.001751\cdot\tau^{22}+0.003828\cdot\tau^{37}.$
Figure 3: The lensing model parameters $a$, $b$, $c$, $p$, and $s$ as functions of $\tau$, depicted for the fluid simulation (dashed lines) and the calibrated parametric model (solid lines). Note that $p(\tau)$ and $s(\tau)$ have no counterparts in the fluid simulation, as they were fixed during a preliminary calibration to better control the fitting of the other parameters. The overall agreement between the two models highlights the accuracy of the parametric calibration.
@@
$\hat{\Sigma}(\hat{R})$ $=\frac{1}{2\hat{R}}\frac{\partial}{\partial\hat{R}}\Bigg{[}\hat{R}\cdot\hat{\alpha}(\hat{R})\Bigg{]}$
@@ 3.2 The density profile model
Although analytic, the equation for the surface density (Eq. 3.1) is already quite complex. Instead of attempting to derive an analytic density profile from the inversion of Eq. 3.1, we propose an alternative density profile model. This model introduces a core in a manner analogous to refs. [1, 90], but includes two additional parameters, which significantly enhance its ability to model density profiles in the deep core-collapse regime ($\tau\gtrsim 0.7$). The new density profile is given by:
@@
$\frac{r_{s}}{r_{s,0}}$ $=1+2.797\cdot\tau^{0.6}-2.687\cdot\tau+3.421\cdot\tau^{3}+0.06970\cdot\tau^{10}+2.027\cdot\tau^{17}+0.03141\cdot\tau^{80}-0.02225\cdot x^{83},$
$\frac{r_{c}}{r_{s,0}}$ $=6.046\cdot\tau^{0.6}-6.957\cdot\tau^{0.8}+1.486\cdot\tau^{2}-0.6234\cdot\tau^{5}+0.2096\cdot\tau^{12}-0.0338\cdot\tau^{27}+5.026\cdot 10^{-5}\cdot\tau^{80},$
$\beta$ $=1-0.4171\cdot\tau^{0.6}+0.2024\cdot\tau^{2}+0.9345\cdot\tau^{13}-2.110\cdot\tau^{26}+2.008\cdot\tau^{39}-0.5762\cdot\tau^{66}+0.2962\cdot\tau^{71}$
$\gamma$ $=2.547-2.756\cdot\tau+6.118\cdot\tau^{2}-7.289\cdot\tau^{3}+3.419\cdot\tau^{4}+\frac{1}{\ln(0.001)}(-1.547)\ln(\tau^{1.330}+0.001)$
$\quad+0.1\cdot\left(\arctan(3(\tau-0.5))-\arctan(-1.5)\right).$
In this model, the parameter $\gamma$ controls the inner density slope when $r_{c}$ is much smaller than $r_{s}$. In the outer region, the profile declines as $r^{-3}$, similar to the NFW profile, while the parameter $\alpha$ allows for control over the shape in the intermediate regions.
In Fig. 2, we present the performance of this model in fitting the fluid simulation results at representative $\tau$ values. We achieve excellent fitting quality across all gravothermal phases in the range $1.080<\tau<$ 1.08. However, for $\tau\gtrsim 1.08$, the halo undergoes a runaway collapse, and the density profile approaches a power-law form in the LMFP regime. This abrupt transition is physical and cannot be smoothly captured by fitting functions. In Section 4, we conduct a detailed study of core-collapsed halos in this regime, where we demonstrate that the PSIDM profile still provides an excellent fit.
@@ 4.1 Isolated halos
Figure 5: Critical curves and caustics for a spherical, isolated SIDM halo. From left to right, the results correspond to progressively larger gravothermal phases with $\tau=0.001$, $0.8$, $1.0$, and $1.05$. Solid curves denote predictions from the parametric model, while dashed curves represent numerical results from the fluid simulation. Both critical curves (red) and caustics (blue) are displayed within the same panels, keeping in mind that they lie on different planes: the lens and the source. Tangential critical curves and caustics are highlighted with darker lines, while radial critical curves and caustics are shown with lighter lines for clarity.
@@
Figure 6: Critical curves and caustics for an isolated SIDM halo with a constant ellipticity $e=0.6$ (or $b/a=0.8$). From left to right, the results correspond to progressively larger gravothermal phases with $\tau=0.001$, $0.8$, $1.0$, and $1.05$. Solid curves denote predictions from the parametric model, while dashed curves represent numerical results from the fluid simulation. Tangential critical curves and caustics are highlighted with darker lines, while radial critical curves and caustics are shown with lighter lines for clarity. The lower panel provides a close-up illustration of the tangential caustics.
@@
Figure 5 compares the model-predicted and numerically simulated critical curves (red) and caustics (blue) for a spherically symmetric halo at $z_{l}=0.439$ with a source plane at $z_{s}=3$. Given our setup, the surface densities within the range $0.019\lesssim\tau\lesssim 0.45$ remain below $\Sigma_{\rm cr}$ for $r>0.1r_{s,0}$. Therefore, we present results for $\tau=0.001$, $0.8$, $1.0$, and $1.05$, from left to right. Our findings show excellent agreement between the two predictions. A discrepancy of $7\%$ in the Einstein radius is observed only at $\tau=0.8$. The tangential caustics vanish consistently in all cases, as a direct consequence of the assumed spherical symmetry.
@@
Importantly, non-zero ellipticity plays a critical role in lensing studies, as it introduces non-vanishing tangential caustics, enabling the formation of four-image lensing events. In SIDM halos, ellipticity is influenced by the thermalization of the inner regions, which reduces ellipticity compared to the CDM case in a radius-dependent manner.
@@
Exploring the effects of a radius-dependent ellipticity is beyond the scope of this work. Instead, we introduce a constant ellipticity manually and numerically compute the corresponding lensing predictions. Figure 6 presents the results for an ellipticity of $e=0.6$ (or $b/a=0.8$). Compared with the spherical case, we observe the emergence of diamond-shaped tangential caustics in the inner regions. Aside from this, the other curves are distorted in shape while maintaining approximately the same area. To better visualize the tangential caustics, we provide zoomed-in views of them in the bottom panels. Comparing the panels at different $\tau$ values, we find that core-collapsing halos, represented by the last three panels, exhibit larger tangential GGSL cross sections than the leftmost panel, which closely resembles the NFW case.
@@
Right panel: The Einstein radius ($\theta_{E}$) as a function of the halo mass ($M$) at different $\tau$ values as color-coded in the bar. The region with $\theta_{E}<1\leavevmode\nobreak$arcsec is shaded in gray, suggesting no detectable signatures. We mark the minimum detectable halo mass in the NFW case ($\tau=0$) and the core collapsed case ($\tau=1.08$) with labeled crosses, revealing that SIDM can reduce this threshold by more than one order of magnitude, from $10^{14.0}M_{\odot}$ to $10^{12.3}M_{\odot}$.
@@ 4.2 Subhalos in a host halo
In this section, we evaluate the effectiveness of a parametric lensing model in describing the critical curves and caustics of subscale lenses. We focus on velocity-dependent SIDM models, which decouple the effective SIDM cross-sections of the host halo and the subhalo. This decoupling allows the host halo to remain unaffected by SIDM effects while enabling the subhalos to evolve into deep core-collapse phases. For clarity, we assume the host halo follows an NFW profile (i.e., $\tau=0$), and we explore the lensing features of subhalos at representative gravothermal evolution phases and varying distances from the host halo. Table 1 summarizes the parameters of the main and host halos used in this study. The main halo adopts the same NFW parameters as the isolated halo discussed in the previous section. The lens redshift ($z_{l}=0.439$) and source redshift ($z_{s}=3$) are also kept consistent with the previous setup. To obtain the Einstein radius $\theta_{E}$ of the host halo, we first identify the radius of the outermost tangential critical curve from the results in figure 5, and use this radius as the Einstein radius of the main halo.
@@ 5 Lensing signatures of core collapsed halos in the self-similar regime
The middle panel presents the effective heat conductivity, defined as $\kappa_{\rm eff}=1/(\kappa_{\rm LMFP}^{-1}+\kappa_{\rm SMFP}^{-1})$, alongside the LMFP (blue) and SMFP (red) ones, for the final snapshot at $\tau=1.08$. This plot demonstrates that the outer halo remains in the LMFP regime even during the late stages of core collapse. Notably, the transition between the two regimes depends on the SIDM model. For the $\sigma/m=5\leavevmode\nobreak\ \rm cm^{2}/g$ case (solid lines), which we adopted in the fluid simulation, the transition occurs at $r/r_{s}\approx 0.02$. Assuming universality and hydrostatic equilibrium, we can extend this result to other SIDM models by holding the simulated halo properties fixed. For instance, for $\sigma/m=100\leavevmode\nobreak\ \rm cm^{2}/g$ (dashed lines), the transition shifts to $r/r_{s}\approx 0.2$, an order of magnitude larger.
@@
To assess the impact of SMFP-induced evolution on the lensing features of a core-collapsed halo, we compare predictions from the simulated $\tau=1.08$ case with two idealized power-law models in figure 12. Both models assume an inner logarithmic density slope of $-2.2$; however, in the power-law-2.2 case (red), the outer region is matched to $\rho_{\rm PSIDM}$, while the Dutra+24 [74] case (dashed blue) follows a generalized NFW profile. We find that the $\tau=1.08$ case introduces only minor differences at the percent level in the radial positions of both the critical curves and caustics, while the two power-law models produce nearly identical results.

### v2

@@ Abstract
We present a model for analytically calculating gravitational lensing by self-interacting dark matter (SIDM) halos. Leveraging the universal behavior of SIDM halos during gravothermal evolution, we calibrate the lensing potential using a fluid simulation, normalizing the evolution time to align with established scenarios. From this potential, we derive explicit equations for the deflection angle and surface density profile, quantifying their deviations from numerical results. Our model builds on the parametric approach of ref. [1], providing refinements in the deep core-collapse regime and enabling more comprehensive lensing studies. We explore characteristic lensing features, including critical curves and caustics, for SIDM halos in isolation and within a main halo, tracking their evolution through the gravothermal phase. We also examine signatures in the self-similar regime of core collapsed halos and highlight the role of baryonic effects in realistic halos. The application of our model extends to generic halos, whose profiles fit one or a superposition of our parametric forms. We make our implementation publicly available on GitHub to support further research.
@@ 1 Introduction
[…]
@@ 2 Universality and simulation data for model calibration
We present a visual representation of the universal gravothermal evolution in figure 1, which can be interpreted as a general solution for a gravothermal system: Given a halo and an SIDM model, the evolution can be derived by specifying the corresponding $\rho_{s,0}$, $r_{s,0}$, and $t_{c}$. The density profile is normalized by the NFW scale parameters at $\tau=0$, with its evolution scaled by $t_{c}$. The corresponding logarithmic density slope, $d\ln(\rho/\rho_{s,0})/d\ln(r/r_{s,0})$, is shown in the right panel. At $\tau=0$, the density profile follows the NFW form, characterized by an inner logarithmic slope of 1-1. By $\tau\approx 1.1$, the density transitions to a power-law profile, $r^{-2.2}$, featuring a core-collapsed halo in the LMFP regime [87, 40]. Dotted and dashed lines are used to track the evolution of the inner halo density, $\rho_{c}$, at $r/r_{s,0}=0.01$, and the core size, $r_{c}$, within the parametric model. Following a sharp decline at $\tau\gtrsim 0$, $\rho_{c}$ reaches its minimum near $\tau\lesssim 0.2$, after which it stabilizes into a quasi-stable core that contracts slowly. As $\tau$ approaches unity, core contraction accelerates significantly, indicating the onset of runaway collapse.
@@
Figure 2: Comparison of SIDM halo density profiles from simulations and model predictions at representative $\tau$ values. The N-body simulation results from ref. [1] were used to calibrate the $\beta 4$ model (orange), while the fluid simulation based on the program of refs. [69, 83] was used to calibrate the PSIDM model (blue) in this work. As discussed in the main text, differences between the previous and new models are minimized for $\tau<0.7$, beyond which the fluid simulation results are expected to provide greater accuracy, particularly in the inner regions. The lower panels show the relative errors of each model with respect to the fluid simulation results.
@@ 3.1 The lensing model
$\Psi(R)=\frac{2}{\Sigma_{\mathrm{cr}}}\int_{0}^{R}s\Sigma(s)\ln\left(\frac{R}{s}\right)ds,$ (3.1)
@@
[…]
Figure 3: The lensing model parameters $a$, $b$, $c$, $p$, and $s$ as functions of $\tau$, depicted for the fluid simulation (dashed lines) and the calibrated parametric model (solid lines). Note that $p(\tau)$ and $s(\tau)$ have no counterparts in the fluid simulation, as they were fixed during a preliminary calibration to better control the fitting of the other parameters. In all cases, the relative differences of the parametric model to the fluid results are consistently below $10^{-4}$.
@@
[…]
@@ 3.2 The density profile model
Although analytic, the equation for the surface density (Eq. 3.1) is already quite complex. Instead of attempting to derive an analytic density profile from the inversion of Eq. 3.1, we propose an alternative density profile model. This model introduces a core in a manner analogous to refs. [1, 88], but includes two additional parameters, which significantly enhance its ability to model density profiles in the deep core-collapse regime ($\tau\gtrsim 0.7$). The new density profile, which we refer to as the PSIDM profile (short for Parametric SIDM), is given by:
@@
[…]
In this model, the parameter $\gamma$ controls the inner density slope when $r_{c}$ is much smaller than $r_{s}$. In the outer region, the profile declines as $r^{-3}$, similar to the NFW profile, while the parameter $\beta$ allows for control over the shape in the intermediate regions.
In Fig. 2, we present the performance of this model in fitting the fluid simulation results at representative $\tau$ values. We achieve excellent fitting quality across all gravothermal phases in the range $0<\tau<1.08$. However, for $\tau\gtrsim 1.08$, the halo undergoes a runaway collapse, and the density profile approaches a power-law form in the LMFP regime. This abrupt transition is physical and cannot be smoothly captured by fitting functions. In Section 4, we conduct a detailed study of core-collapsed halos in this regime, where we demonstrate that the PSIDM profile still provides an excellent fit.
@@ 4.1 Isolated halos
Figure 5: Critical curves and caustics for a spherical, isolated SIDM halo. From left to right, the results correspond to progressively larger gravothermal phases with $\tau=0.001$, $0.8$, $1.0$, and $1.05$. Solid curves denote predictions from the parametric model, while dashed curves represent numerical results from the fluid simulation. Both critical curves (red) and caustics (blue) are displayed within the same panels, keeping in mind that they lie on different planes: the lens and the source. Tangential critical curves and caustics are highlighted with darker lines, while radial critical curves and caustics are shown with lighter lines for clarity. $|\Delta\theta_{E}|$ denotes the absolute difference between the effective Einstein radii obtained from the parametric model and the fluid simulation.
@@
Figure 6: Critical curves and caustics for an isolated SIDM halo with a constant ellipticity $e=0.6$ (or $b/a=0.8$). From left to right, the results correspond to progressively larger gravothermal phases with $\tau=0.001$, $0.8$, $1.0$, and $1.05$. Solid curves denote predictions from the parametric model, while dashed curves represent numerical results from the fluid simulation. Tangential critical curves and caustics are highlighted with darker lines, while radial critical curves and caustics are shown with lighter lines for clarity. The lower panel provides a close-up illustration of the tangential caustics. $|\Delta\theta_{E}|$ is defined consistently with Figure 5 and $A_{T}$ refers to the area enclosed by the tangential caustic.
@@
Figure 5 compares the model-predicted and numerically simulated critical curves (red) and caustics (blue) for a spherically symmetric halo at $z_{l}=0.439$ with a source plane at $z_{s}=3$. Given our setup, the surface densities within the range $0.019\lesssim\tau\lesssim 0.45$ remain below $\Sigma_{\rm cr}$ for $r>0.1r_{s,0}$. Therefore, we present results for $\tau=0.001$, $0.8$, $1.0$, and $1.05$, from left to right. Our findings show excellent agreement between the two predictions. The effective Einstein radius is defined as $\theta_{E}=\sqrt{A_{E}/\pi}$, where $A_{E}$ represents the area enclosed by the tangential critical curve. A discrepancy of 7% in the Einstein radius is observed only at $\tau=0.8$. Due to the assumed spherical symmetry, the tangential caustics in the source plane degenerate into single points at the origin, rather than forming extended curves. These correspond to circular critical curves in the image plane.
@@
Importantly, non-zero ellipticity plays a critical role in lensing studies, as it introduces non-vanishing tangential caustics, enabling the formation of four-image lensing events. In this context, galaxy-galaxy strong lensing (GGSL) cross section quantifies the area on the source plane within which a background source would be multiply imaged due to the lensing potential of a foreground halo. It is highly sensitive to both the inner mass distribution and ellipticity of the lens. In SIDM halos, ellipticity is influenced by the thermalization of the inner regions, which reduces ellipticity compared to the CDM case in a radius-dependent manner.
@@
Exploring the effects of a radius-dependent ellipticity is beyond the scope of this work. Instead, we introduce a constant ellipticity using the coordinate transformation method of Ref. [75], and numerically compute the resulting lensing predictions. The ellipticity is defined as $e=\sqrt{1-q^{2}}$, where $q=b/a$ is the axis ratio of the projected mass distribution. Figure 6 presents the results for an ellipticity of $e=0.6$ (or $q=b/a=0.8$). Compared with the spherical case, we observe the emergence of diamond-shaped tangential caustics in the inner regions. Aside from this, the other curves are distorted in shape while maintaining approximately the same area. To better visualize the tangential caustics, we provide zoomed-in views of them in the bottom panels. Comparing the panels at different $\tau$ values, we find that core-collapsing halos, represented by the last three panels, exhibit larger tangential GGSL cross sections than the leftmost panel, which closely resembles the NFW case.
@@
Right panel: The Einstein radius ($\theta_{E}$) as a function of the halo mass ($M$) at different $\tau$ values as color-coded in the bar. The region with $\theta_{E}<1\leavevmode\nobreak\$arcsec is shaded in gray, suggesting no detectable signatures. We mark the minimum detectable halo mass in the NFW case ($\tau=0$) and the core collapsed case ($\tau=1.08$) with labeled crosses, revealing that SIDM can reduce this threshold by more than one order of magnitude, from $10^{14.0}M_{\odot}$ to $10^{12.3}M_{\odot}$.
@@ 4.2 Subhalos in a host halo
In this section, we evaluate the effectiveness of a parametric lensing model in describing the critical curves and caustics of subscale lenses. We focus on velocity-dependent SIDM models, which decouple the effective SIDM cross-sections of the host halo and the subhalo. This decoupling allows the host halo to remain unaffected by SIDM effects while enabling the subhalos to evolve into deep core-collapse phases. For clarity, we assume the host halo follows an NFW profile (i.e., $\tau=0$), and we explore the lensing features of subhalos at representative gravothermal evolution phases and varying distances from the host halo. Table 1 summarizes the parameters of the main halo and subhalo used in this study, chosen to enhance strong lensing features and more clearly demonstrate the accuracy of our calibrated model. The main halo adopts the same NFW parameters as the isolated halo discussed in the previous section. The lens redshift ($z_{l}=0.439$) and source redshift ($z_{s}=3$) are also kept consistent with the previous setup. To obtain the Einstein radius $\theta_{E}$ of the host halo, we first identify the radius of the tangential critical curve from the results in figure 5, and use this radius as the Einstein radius of the main halo.
@@ 5 Lensing signatures of core collapsed halos in the self-similar regime
In the conducting fluid model of SIDM halos, self-interactions are modeled through a heat transport equation, where the scattering cross section determines the thermal conductivity $\kappa$. Unlike ordinary fluids, SIDM halos primarily exhibit a long-mean-free-path (LMFP) regime, in which particles can orbit the halo center multiple times after a collision, suppressing the energy transport. This contrasts with the short-mean-free-path (SMFP) regime typical of fluid dynamics. To capture both regimes and their transition during gravothermal evolution, the effective conductivity is defined as $\kappa_{\rm eff}=\left(\kappa_{\rm LMFP}^{-1}+\kappa_{\rm SMFP}^{-1}\right)^{-1}$, providing a smooth interpolation between the LMFP and SMFP regimes [40, 60, 41, 59, 61, 62, 64].
@@
The middle panel presents the distribution of $\kappa m$ for the effective (black), LMFP (blue), and SMFP (red) scenarios, for the final snapshot at $\tau=1.08$. This plot demonstrates that the outer halo remains in the LMFP regime even during the late stages of core collapse. Notably, the transition between the two regimes depends on the SIDM model. For the $\sigma/m=5\leavevmode\nobreak\ \rm cm^{2}/g$ case (solid lines), which we adopted in the fluid simulation, the transition occurs at $r/r_{s}\approx 0.02$. Assuming universality and hydrostatic equilibrium, we can extend this result to other SIDM models by holding the simulated halo properties fixed. For instance, for $\sigma/m=100\leavevmode\nobreak\ \rm cm^{2}/g$ (dashed lines), the transition shifts to $r/r_{s}\approx 0.2$, an order of magnitude larger.
@@
To assess how different descriptions of core-collapsed halos affect lensing features, we compare predictions from the simulated $\tau=1.08$ case with two idealized power-law models. The first model, labeled power-law-2.2, adopts an inner density profile proportional to $r^{-2.2}$, smoothly matched to the outer region of $\rho_{\rm PSIDM}$. The second model, following Dutra et al. [73], uses a generalized NFW profile [89]:
$\rho_{\rm gNFW}=\frac{\rho_{0}}{(r/r_{0})^{\gamma_{1}}\left(1+(r/r_{0})^{2}\right)^{(\gamma_{2}-\gamma_{1})/2}},$ (5.1)
where we set $\gamma_{1}=2.2$ and $\gamma_{2}=3$, allowing $r_{0}$ and $\rho_{0}$ to vary in the fit. Figure 12 compares the density profiles (left), critical curves (middle), and caustics (right) for the fluid simulation at $\tau=1.08$ (orange), power-law-2.2 (red), and Dutra+24 (blue). We find that the $\tau=1.08$ case introduces only minor differences at the percent level in the radial positions of both the critical curves and caustics, while the two power-law models produce nearly identical results.